# Prethermalization in Fermi-Pasta-Ulam-Tsingou chains

### v1

@@ IV.2 Equipartition and Lyapunov times
Figure 5: Equipartition times computed by evolving the first five odd roots (rainbow colors) and 280 random initial states (tones of magenta and orange) at several different energy densities. These data were obtained from spectral entropies computed up to $t{=}10^{8}$, such that the top left equipartition times, which are of $\mathcal{O}(10^{8})$, are likely slightly overestimated by being of the order of the computation time. The equipartition time for $Q_{1}$ (red diamonds) clearly forms an upper bound that the ones computed from random initial states essentially never cross. The equipartition times computed from excited modes consistently do so, albeit slowly as a function of increasing root number.
@@
If equipartition times reach the order of the propagation time used to compute them, namely $\mathcal{O}(10^{8})$, convergence can no longer be achieved for these parameters, which is clear in Fig. 5 by the presence of missing data for some values of energy density. However, the “disappearance” of points does not increase fully monotonically with mode excitation number, with some data displaying what could be described as “erratic behavior”. Since dependence on initial conditions is tantamount to ergodicity breaking, investigating quantities that are proved to be invariant in any ergodic system will shed light into the source of the instabilities seen in Fig. 5. The most important of such quantities is the MLE, $\lambda_{1}$, which indicates the presence of chaotic behavior and is not only invariant with respect to diffeomorphisms but also constant for ergodic dynamical systems [27]. From this exponent one can extract the Lyapunov time, $\tau_{1}=1/\lambda_{1}$, which measures the time a vector tangent to a trajectory takes to align with the maximally chaotic, one-dimensional submanifold in the tangent bundle [39, 40]. Since the Lyapunov time is computed from the divergence of two initially close trajectories, it takes a finite time to converge and it is often useful to observe its evolution as a function of time, as done in Fig. 6.
@@
Figure 6: (a) Time-dependent Lyapunov times computed by evolving the first five odd modes at a fixed energy density of $h{\approx}0.085$ up to $t_{f}{=}10^{9}$. Clearly, at this energy density all modes behave similarly and indistinguishably from the random initial states shown in panel (b), for which the propagation time was set to $t_{f}{=}10^{8}$. (c) By lowering the energy density down to $h{\approx}0.033$, roots start to behave differently as a function of time when compared to the random states used in (d). The differences appear to grow with excitation number, resulting in high-order modes such $Q_{9}$ not converging at all within this propagation interval (which is more than enough for all other initial states).
@@
Panels (a) and (b) of Fig. 6 display the evolution of $\tau_{1}(t)$ extracted from propagating odd roots and random initial conditions for an energy density of $h{\approx}0.085$. Not only is convergence achieved for all initial states employed, but the time-dependent portrait of both typical and atypical initial conditions is identical. Upon decreasing the energy density, however, Fig. 6(c) shows that atypical initial conditions start behaving very differently not only with respect to typical ones, but also among themselves. While the Lyapunov times obtained from $Q_{1}$ and $Q_{3}$ are similar to the ones of the random initial conditions in Fig. 6(d), the data obtained from evolving $Q_{5}$ is seen to start forming a “belly” before convergence. Thus, it shows that $Q_{5}$ lies in a region of phase space that takes longer to align with the maximally chaotic direction, similarly to the trapping times described in [23]. The difference is that here we observe longer trapping times for a fixed energy density by changing only the initial conditions.
@@
Figure 7: (a) Lyapunov times computed by evolving the first five odd roots (rainbow colors) at several different energy densities. Propagation time used was $t_{f}{=}10^{9}$, for which $\tau_{1}$’s computed by evolving $Q_{1}$ converge well for all energy densities used. Other modes, however, did not converge for the whole range of energy densities employed here, requiring longer and unpractical propagation times (see text for explanation). (b) Lyapunov times computed by evolving 10 random initial states at the same energy densities as panel (a). Propagation times in this case were only $t_{f}{=}10^{7}$, which is enough to obtain converged results starting from typical initial conditions. Error bars in all plots are the standard deviation of the corresponding $\tau_{1}(t)$ between $t_{f}$ and $t_{f}/10$.
By evolving even higher excited roots such as $Q_{7}$, Fig. 6(c) shows that the “belly” increases when compared to the one of $Q_{5}$, such that $Q_{7}$ is trapped for a time one order of magnitude longer. Extrapolating to $Q_{9}$ this transient trapping regime, which is a manifestation of prethermalization in an invariant quantity and therefore must disappear for sufficiently long times, ends up being longer than the propagation time and does not converge in our computations. The reason for the longer transient times as we go up from $Q_{1}$ to $Q_{9}$ is that the higher the excitation number, the smaller the subspace of mode space in which dynamics is approximately restricted to during the prethermal regime, as discussed in Sec. III. Indeed, Fig. 3 shows that a large fraction of the energy is still concentrated in multiples of the initial excitation number at equipartition time, which as seen by comparing Figs. 6 and 7 is always longer than Lyapunov time when dealing with atypical initial conditions. Thus, as long as the root is a low lying mode (that is, $Q_{r}$ with $2(r-1)<N$), the trapping time will increase with the excitation number. The breather at $Q_{32}$ marks the boundary between low-mode and high-mode excitations, with $Q_{31}$ being the slowest mode to thermalize for $N{=}63$–so slow that there is no point in attempting to obtain data from it, since this is already computationally hard for $Q_{9}$. It is interesting to note that components of the trajectory that “scan” phase space in order to obtain convergence in $\tau_{1}$ are the low-energy ripples that are not multiples of $r$ for a root $Q_{r}$. Evidently, these ripples might not have enough energy to go very far, requiring longer transient times. The energy is spread among all modes most efficiently when evolving $Q_{1}$, which explains why this state behaves essentially as a random state despite its extremely long equipartition time when compared to the latter.
@@
Figure 8: Time-dependent Kolmogorov-Sinai entropy in an $\alpha$-FPUT chain with $N{=}63$ masses computed from roots (rainbow colors) and from a random state (black). Each panel is computed for initial states with decreasing energy densities, $h$, shown in the left corner.
@@
Fig. 7, which shows the final (converged) Lyapunov times as a function of energy density as obtained from typical and atypical initial conditions, is clear evidence that the system we are dealing with is numerically ergodic: The times are the same, no matter what initial conditions are chosen (although the larger error bars and spread for $h{\approx}10^{-2}$ show that $t_{f}{=}10^{9}$ is barely long enough to reach numerical ergodicity for the smallest values of $h$ used here). This does not mean, as discussed before, that all data for roots converges, with the main reason for failure being the presence of extremely long prethermal regimes as seen in Fig. 6.
@@ IV.3 Kolmogorov-Sinai entropies
Given the imprint prethermalization has in a dynamical invariant such as the Lyapunov time, in Fig. 8 we display the time evolution of one of the most meaningful, yet computationally expensive, invariant quantities in dynamical systems, namely the Kolmogorov-Sinai entropy, $\kappa_{\mathrm{KS}}$. This quantity, just like the Lyapunov time and the associated maximal exponent, displays a clear dip when computed by propagating roots, while for a typical initial condition it approaches its converged value from above. Since $\kappa_{\mathrm{KS}}$ is the sum of all positive Lyapunov exponents, what Fig. 8 shows is that the prethermal regime is present in all phase-space directions, not only the maximal one associated to the MLE. Upon decreasing the energy density we also see a tendency of a larger dip for higher-excited modes, in accordance with our expectation that the prethermal regime is more pronounced for higher excitation numbers. We also note that convergence of Kolmogorov-Sinai entropies to the same final value is likely the strongeast display of numerical ergodicity achievable in simulations.
@@ V Discussion
Traditionally, the Lyapunov time is considered to characterize alignment along the maximally chaotic direction in a decomposition of the tangent space known as Oseledec splitting, which is covariant with respect to the Hamiltonian flow [40]. Evidently, such decomposition is meaningless in the case of a near-regular trajectory, and it is expected that the transient regime of $\lambda_{1}(t)$ will be different depending on how close the initial state is to a near-integrable region. However, the Lyapunov time itself is an invariant quantity in ergodic systems, such that at the end all initial states must provide the same value for the converged MLE (that is, the asymptotic limit of $\lambda_{1}(t))$. This was verified in Figs. 6 and 7. Thus, the time needed to escape a region of near-integrability is more related to the equipartition time than to Lyapunov time, since we expect mode space to be fully covered only once a trajectory starting in a root becomes sufficiently chaotic. The Lyapunov time itself then carries absolutely no information on prethermalization, although its time-evolution certainly does. This leads to a reinterpretation of $\tau_{1}$ as the time a typical initial condition takes to align with the maximally chaotic covariant direction in tangent space, since atypical ones might not align with it at all. This results in the Lyapunov time being always shorter than the equipartition time of a root, as seen by plotting them together in Fig. 9.
@@
It would be interesting to attempt a computation of the Lyapunov time of a submanifold, i.e. calculate $\tau_{1}$ starting from a nonthermal mode, e.g. $Q_{2}$, and see how this exponent relates to the true $\tau_{1}$ obtained from a thermal initial condition. Unfortunately this type of numerical investigation is hard, if not impossible, to carry on. This is due to the fact that the bushes associated to nonthermal modes in $\alpha$-FPUT chains are unstable and will not be preserved in long-time evolution no matter which numerical integrator is chosen [26]. Indeed, if the root chosen is nonthermal, by tracking the time-evolution of mode energies one sees that the initial excitation leaves the bush due to numerical errors and eventually covers the whole of mode space. This instability might be responsible for the observation that the evolution of nonthermal modes results in approximately the same Lyapunov times as thermal ones [23], which is surprising given that bush dynamics takes place only in a submanifold of mode space. Nevertheless, it is unclear how ergodicity in mode space relates to that of phase space, such that it is possible that a partial covering of mode space by a large enough bush is enough to resolve the “true” MLE. Besides, it might even happen that dynamics in phase space is ergodic while that in mode space is not, since ergodicity always depends on the observables chosen [41] unless it is tracked by means of invariant quantities [42]. At present, the authors are unaware of a special type of integrator that is capable of preserving bushes of excitations, such that studies regarding dynamical properties of chaotic subspaces cannot be performed in a numerically meaningful fashion. A consequence of this is that the energy in Fig. 1(a), which is correctly localized in the bush associated to $Q_{3}$, will eventually spread to all modes for long enough times. Nevertheless, the way “equipartition” takes place in this case is pathological and clearly traceable to numerical errors, as can be seen in Appendix A.
@@ VI Conclusion
We have extended previous studies on the cubic Fermi-Pasta-Ulam-Tsingou ($\alpha$-FPUT) model by investigating the consequences of evolving higher-excited normal modes instead of the fundamental. The evolution of such atypical initial conditions was also compared to that obtained from typical (random) ones. We provided an explicit and simple way of predicting mode excitation sequences when exciting only a single normal mode (the root), showing that higher-excited roots take longer to cover mode space, while the fundamental is the fastest energy spreader. Thorough numerical investigations based on computing Lyapunov exponents and Kolmogorov-Sinai entropies show that even invariant quantities carry imprints of the atypical initial states used to compute them in their transient regime, despite converging to the correct invariant value for long times. The prethermal dynamics taking place before convergence is confirmed to be a consequence of some select roots lying close to stationary states also in the $\alpha$-FPUT system, and therefore being trapped in near-regular regions. The escape from such regions is what puts an end to the paradigmatic metastable plateaus in mode energies, whose end marks the moment invariant quantities computed from roots start to converge.
@@ Acknowledgements.
This research was supported by the Institute for Basic Science through Project Code (No. IBS-R024-D1).
@@ Appendix A Numerical instability of bushes
Figure 10: Examples of error-induced equipartition due to numerical instability of bushes. (a) Same as Fig. 1(a) but for longer times. (b) The evolution of $Q_{2}$, which as all even roots is nonthermal, in an $N{=}11$ chain. In both panels, states that are proven to be nonthermal end up reaching equipartition due to numerical errors. Panels (a) and (b) use different symplectic integrators (second and fourth orders), showing that the phenomenon happens independently of the solver, although it happens at different times depending on the time step (attesting for its numerical source).
The bushes associated with nonthermal roots in $\alpha$-FPUT chains appear to be unstable as they evolve in time. The main consequence of this instability is that energy, which should be only shared among modes in the bush and never leave for any finite time, will end up exciting other modes due to numerical errors. In Fig. 10 we provide two examples of this behavior when evolving provenly nonthermal modes: The roots $Q_{3}$ and $Q_{2}$ in a $N=11$ chain. The first case is essentially a continuation of Fig. 1 for much longer times, until one can see an anomalous and simultaneous jump in the energy of modes that should not be excited starting from $Q_{3}$; The root $Q_{2}$, on the other hand, is even and therefore always nonthermal, undergoing a very similar pathological jump and reaching equipartition in Fig. 10 exclusively due to numerical errors.

### v2

@@ IV.2 Equipartition and Lyapunov times
Figure 5: Equipartition times computed by evolving the first five odd roots (rainbow colors) and 280 random initial states (tones of magenta and orange) at several different energy densities in a chain with $N{=}63$ and $\alpha{=}1/4$. These data were obtained from spectral entropies computed up to $t{=}10^{8}$, such that the top left equipartition times, which are of $\mathcal{O}(10^{8})$, are likely slightly overestimated by being of the order of the computation time. The equipartition time for $Q_{1}$ (red diamonds) clearly forms an upper bound that the ones computed from random initial states essentially never cross. The equipartition times computed from excited modes consistently do so, albeit slowly as a function of increasing root number.
@@
If equipartition times reach the order of the propagation time used to compute them, namely $\mathcal{O}(10^{8})$, convergence can no longer be achieved for these parameters, which is clear in Fig. 5 by the presence of missing data for some values of energy density. However, the “disappearance” of points does not increase fully monotonically with mode excitation number, with some data displaying what could be described as “erratic behavior”. Since dependence on initial conditions is tantamount to ergodicity breaking, investigating quantities that are proved to be invariant in any ergodic system will shed light into the source of the instabilities seen in Fig. 5. The most important of such quantities is the MLE, $\lambda_{1}$, which indicates the presence of chaotic behavior and is not only invariant with respect to diffeomorphisms but also constant for ergodic dynamical systems [28]. From this exponent one can extract the Lyapunov time, $\tau_{1}=1/\lambda_{1}$, which measures the time a vector tangent to a trajectory takes to align with the maximally chaotic, one-dimensional submanifold in the tangent bundle [33, 44, 45]. Since the Lyapunov time is computed from the divergence of two initially close trajectories, it takes a finite time to converge and it is often useful to observe its evolution as a function of time, as done in Fig. 6.
@@
Figure 6: (a) Time-dependent Lyapunov times computed by evolving the first five odd modes of a chain with $N{=}63$ and $\alpha{=}1/4$ at a fixed energy density $h{\approx}0.085$ up to $t_{f}{=}10^{9}$. Clearly, at this energy density all modes behave similarly and indistinguishably from the random initial states shown in panel (b), for which the propagation time was set to $t_{f}{=}10^{8}$. (c) By lowering the energy density down to $h{\approx}0.033$, roots start to behave differently as a function of time when compared to the random states used in (d). The differences appear to grow with excitation number, resulting in high-order modes such $Q_{9}$ not converging at all within this propagation interval (which is more than enough for all other initial states).
@@
Panels (a) and (b) of Fig. 6 display the evolution of $\tau_{1}(t)$ extracted from propagating odd roots and random initial conditions for an energy density of $h{\approx}0.085$. Not only is convergence achieved for all initial states employed, but the time-dependent portrait of both typical and atypical initial conditions is identical. Upon decreasing the energy density, however, Fig. 6(c) shows that atypical initial conditions start behaving very differently not only with respect to typical ones, but also among themselves. While the Lyapunov times obtained from $Q_{1}$ and $Q_{3}$ are similar to the ones of the random initial conditions in Fig. 6(d), the data obtained from evolving $Q_{5}$ is seen to start forming a “belly” before convergence. Thus, it shows that $Q_{5}$ lies in a region of phase space that takes longer to align with the maximally chaotic direction, similarly to the trapping times described in [24]. The difference here is that we observe longer trapping times at fixed energy density as the root number increases.
@@
Figure 7: (a) Lyapunov times computed by evolving the first five odd roots (rainbow colors) of a chain with $N{=}63$ and $\alpha{=}1/4$ at several different energy densities. Propagation time used was $t_{f}{=}10^{9}$, for which $\tau_{1}$’s computed by evolving $Q_{1}$ converge well for all energy densities used. Other modes, however, did not converge for the whole range of energy densities employed here, requiring longer and unpractical propagation times (see text for explanation). (b) Lyapunov times computed by evolving 10 random initial states at the same energy densities as panel (a). Propagation times in this case were only $t_{f}{=}10^{7}$, which is enough to obtain converged results starting from typical initial conditions. Error bars in all plots are the standard deviation of the corresponding $\tau_{1}(t)$ between $t_{f}$ and $t_{f}/10$.
@@
Figure 8: Time-dependent Kolmogorov-Sinai entropy in an $\alpha$-FPUT chain with $N{=}63$ and $\alpha{=}1/4$ computed from roots (rainbow colors) and from a random state (black). Each panel is computed for initial states with decreasing energy densities, $h$, shown in the left corner.
By evolving even higher excited roots such as $Q_{7}$, Fig. 6(c) shows that the “belly” increases when compared to the one of $Q_{5}$, such that $Q_{7}$ is trapped for a time one order of magnitude longer. Extrapolating to $Q_{9}$ this transient trapping regime, which is a manifestation of prethermalization in an invariant quantity and therefore must disappear for sufficiently long times, ends up being longer than the propagation time and does not converge in our computations. The reason for the longer transient times as we go up from $Q_{1}$ to $Q_{9}$ is that the higher the excitation number, the smaller the subspace of mode space to which dynamics is approximately restricted during the prethermal regime and the higher the mode energies, as discussed in Sec. III.5. Indeed, Fig. 3 shows that a large fraction of the energy is still concentrated in multiples of the initial excitation number at equipartition time, which as seen by comparing Figs. 6 and 7 is always longer than Lyapunov time when dealing with atypical initial conditions. Thus, the trapping time will increase with the excitation number, which is also a consequence of higher-excited modes lying closer to q-breathers.
@@
Fig. 7, which shows the final (converged) Lyapunov times as a function of energy density for typical and atypical initial conditions, is clear evidence that the system we are dealing with is numerically ergodic: The times are the same, no matter what initial conditions are chosen (although the larger error bars and spread for $h{\approx}10^{-2}$ show that $t_{f}{=}10^{9}$ is barely long enough to reach numerical ergodicity for the smallest values of $h$ used here). This does not mean, as discussed before, that all data for roots converges, with the main reason for failure being the presence of extremely long prethermal regimes as seen in Fig. 6.
@@ IV.3 Kolmogorov-Sinai entropies
Given the imprint prethermalization has in a dynamical invariant such as the Lyapunov time, in Fig. 8 we display the time evolution of one of the most meaningful, yet computationally expensive, invariant quantities in dynamical systems, namely the Kolmogorov-Sinai entropy, $\kappa_{\mathrm{KS}}$. This quantity, just like the Lyapunov time and the associated maximal exponent, displays a clear dip when computed by propagating roots, while for a typical initial condition it approaches its converged value from above. Since $\kappa_{\mathrm{KS}}$ is the sum of all positive Lyapunov exponents, what Fig. 8 shows is that the prethermal regime is present in all phase-space directions, not only the maximal one associated to the MLE. Upon decreasing the energy density we also see a tendency of a larger dip for higher-excited modes, in accordance with our expectation that the prethermal regime is more pronounced for higher excitation numbers. It is worth noting that the convergence of Kolmogorov–Sinai entropies to the same asymptotic value likely constitutes the most compelling numerical evidence of ergodicity attainable in simulations.
@@ V Discussion
Traditionally, the Lyapunov time is considered to characterize alignment along the maximally chaotic direction in a decomposition of the tangent space known as Oseledec splitting, which is covariant with respect to the Hamiltonian flow [33, 45]. Evidently, such decomposition is meaningless in the case of a near-regular trajectory, and it is expected that the transient regime of $\lambda_{1}(t)$ will be different depending on how close the initial state is to a near-integrable region. However, the Lyapunov time itself is an invariant quantity in ergodic systems, such that at the end all initial states must provide the same value for the converged MLE (that is, the asymptotic limit of $\lambda_{1}(t))$. This was verified in Figs. 6 and 7. Thus, the time needed to escape a region of near-integrability is more related to the equipartition time than to Lyapunov time, since we expect mode space to be fully covered only once a trajectory starting in a root becomes sufficiently chaotic. The Lyapunov time itself then carries absolutely no information on prethermalization, although its time-evolution certainly does. This leads to a reinterpretation of $\tau_{1}$ as the time a typical initial condition takes to align with the maximally chaotic covariant direction in tangent space, since atypical ones might not align with it at all. This results in the Lyapunov time being always shorter than the equipartition time of a root, as seen by plotting them together in Fig. 9.
@@
It would be interesting to attempt a computation of the Lyapunov time of a submanifold, i.e., calculate $\tau_{1}$ starting from a nonthermal mode, e.g., $Q_{2}$, and see how this exponent relates to the true $\tau_{1}$ obtained from a thermal initial condition. Unfortunately this type of numerical investigation is hard, if not impossible, to carry on. This is due to the fact that the bushes associated to nonthermal modes in $\alpha$-FPUT chains are unstable and will not be preserved in long-time evolution no matter which numerical integrator is chosen [27]. Indeed, if the chosen root is nonthermal, the time evolution of the mode energies reveals that the initial excitation eventually escapes the bush and spreads throughout mode space, as numerical integration does not preserve the discrete symmetries of bushes. This might be behind the observation that the evolution of nonthermal modes results in approximately the same Lyapunov times as thermal ones [24], which is surprising given that bush dynamics takes place only in a submanifold of mode space. Nevertheless, it is unclear how ergodicity in mode space relates to that of phase space, such that it is possible that a partial covering of mode space by a large enough bush is enough to resolve the “true” MLE. Besides, it might even happen that dynamics in phase space is ergodic while that in mode space is not, since ergodicity always depends on the observables chosen [46] unless it is tracked by means of invariant quantities [47]. At present, the authors are unaware of a special type of integrator that is capable of preserving bushes of excitations, such that studies regarding dynamical properties of chaotic subspaces cannot be performed in a numerically meaningful fashion. A consequence of this is that the energy in Fig. 1(a), which is correctly localized in the bush associated to $Q_{3}$, will eventually spread to all modes for long enough times. Nevertheless, the way “equipartition” takes place in this case is pathological and clearly traceable to numerical errors, as can be seen in Appendix A.
@@ VI Conclusion
We have extended previous studies on the cubic Fermi-Pasta-Ulam-Tsingou ($\alpha$-FPUT) model by investigating the consequences of evolving higher-excited normal modes instead of the fundamental. The evolution of such atypical initial conditions was also compared to that obtained from typical (random) ones. We provided an explicit and simple way of ordering excited modes perturbatively in $\alpha$ when starting from a single normal mode (the root), and showed that higher-excited roots take longer to cover mode space. Thorough numerical investigations based on computing Lyapunov exponents and Kolmogorov-Sinai entropies showed that even invariant quantities carry imprints of the atypical initial states used to compute them in their transient regime, despite converging to the correct invariant value for long times. The prethermal dynamics taking place before convergence is confirmed to be a consequence of some select roots lying close to stationary states also in the $\alpha$-FPUT system, and therefore being trapped in near-regular regions. The escape from such regions is what puts an end to the paradigmatic metastable plateaus in mode energies, whose end marks the moment invariant quantities computed from roots start to converge.
@@ Acknowledgements.
This research was supported by the Institute for Basic Science through project IBS-R024-D1.
@@ Appendix A Numerical instability of bushes
Figure 10: Examples of error-induced equipartition due to numerical instability of bushes in a chain with $N{=}11$, $h{\approx}0.16$ and $\alpha{=}1/4$. (a) Same as Fig. 1(a) but for longer times. (b) The evolution of $Q_{2}$, which as all even roots is nonthermal. In both panels, states that are proven to be nonthermal end up reaching equipartition due to numerical errors. Panels (a) and (b) use different symplectic integrators (second and fourth orders), showing that the phenomenon happens independently of the solver, although it happens at different times depending on the time step (attesting for its numerical source).
The bushes associated with nonthermal roots in $\alpha$-FPUT chains obey discrete symmetries that are not respected by the integrators, being therefore unstable in numerical simulations. The main consequence of this instability is that energy, which should only be shared among modes in the bush and never leave for any finite time, ends up exciting other modes due to numerical errors. In Fig. 10 we provide two examples of this behavior when evolving provenly nonthermal modes: the roots $Q_{3}$ and $Q_{2}$ in a $N=11$ chain. The first case is essentially a continuation of Fig. 1 for much longer times, until one can see an anomalous and simultaneous jump in the energy of modes that should not be excited starting from $Q_{3}$; the root $Q_{2}$, on the other hand, is even and therefore always nonthermal, undergoing a very similar pathological jump and reaching equipartition in Fig. 10 exclusively due to numerical errors.